Give in set notation the family described:
All lines containing(2,5)

Answer by Alan3354(69443)

Give in set notation the family described:
All lines containing(2,5)
---------
y-5 = m*(x-2) for all real values of m
m = -k is the same as m = k (k is any real number)
--> y-5 = |m|*(x-2)
m = 0 to infinity